pure acid is to be added to a 10% acid solution to obtain a 24 L OK 40% acid solution. What amount should be used?
Answer by ankor@dixie-net.com(22740)

pure acid is to be added to a 10% acid solution to obtain a 24 L OK 40% acid solution.
What amount should be used?
:
let a = amt of pure acid required to accomplish this
The total amt is to be 24 L, therefore
(24-a) = amt of 10% acid required
:
using the decimal equiv of per cent
1a + .10(24-a) = .40(24)
a + 2.4 - .10a = 9.6
a - .10a = 9.6 - 2.4
.9a = 7.2
a = 7.2/.9
a = 8 liters of pure acid required